OpenStudy (anonymous):

The chart shows the weight of a puppy. How much weight did the puppy gain between Week 1 and Week 4?

OpenStudy (anonymous):

heres the chart

OpenStudy (anonymous):

subtract the two quantities.... week4: \(\large 3\frac{2}{8} \) week1: \(\large 2\frac{3}{8} \) so amount gain = week4 - week1 = \(\large 3\frac{2}{8} -2\frac{3}{8} =\) ????

OpenStudy (anonymous):

so it would be 1 1/0

OpenStudy (anonymous):

no.... you'll need to "borrow" a 1 from the 3 and so now you'll have: \(\large 3\frac{2}{8} -2\frac{3}{8} =2\frac{10}{8} -2\frac{3}{8}= \)???

OpenStudy (anonymous):

it would be 7

OpenStudy (anonymous):

7/8... you're subtracting fractions so don't forget the denominator....